OpenStudy (anonymous):

graph the line 5x+2y=8

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study..........................

OpenStudy (anonymous):

Anywho, what he said^ but to help you with your question :) You're going to want to convert this equation into slope-intercept form. 1) isolate the 'y' variable :) 2y = -5x + 8 (whatever you do to one side you MUST do to the other) 2) divide each side by two y = (-5x + 8)/(2) y = -5/2x + 4 So slope-intercept form gives you the 'y' intercept and the slope. y = mx+b where m is the slope and b is the y-intercept so the slope is -5/2 and the y-intercept is (0, 4)

OpenStudy (anonymous):

so slope is \[\frac{ rise }{ run }\] using that, since it's -5, we want to go DOWN from (0, 4) f coordinates, and over to the LEFT 2 coordinates.

OpenStudy (anonymous):

or, you could use a graphing calculator and graph the equation y = -5/2x + 4 but I recommend what I'm trying to do. But I'm not sure how to show you what I'm doing